If you wish to lose weight, you may rub a balloon on your head. If you charge the balloon to a charge of q = -1.23µC, how much weight did you lose? The mass of an electron is 9.11 × 10-31 kg.
Answer #1

using quantization of charge

n = q / e = 1.23*10^-6 / ( 1.6*10^-19)

n = 7.6875*10^12

weight of n electron

W = n * m g

W = 7.6875*10^12* 9.8* 9.11*10^-31

W = 6.863*10^-17 N

==========

Comment in case any doubt, will reply for sure.. Goodluck